# Rethinking Prompt Optimizers: From Prompt Merits to Optimization

### v2

@@ 4.2 Main Results
PattoPO 83.33 68.52 83.12 54.35 83.46 74.56
Tulu2-7b-dpo Training Free 49.58 34.30 32.93 40.96 72.60 46.22
PattoPO 55.05 38.14 35.18 43.25 73.60 49.04
LLaMA2-7b-chat-hf Training Free 36.91 30.55 20.91 39.88 53.10 36.27
PattoPO 39.86 31.74 25.25 41.97 55.33 38.83
LLaMA3-8b-instruct Training Free 52.36 50.77 58.25 42.80 71.69 55.17
PattoPO 55.01 51.88 59.59 45.52 76.66 57.73
Table 3: Comparison between training-free merit-guided prompts and MePO. Detailed BBH results are in Appx.D.3.
Downward Compatibility FIPO shows performance degradation compared to the inference model optimizer on LLaMA2-7b (-0.5% on ARC-Easy, -0.08% on ARC-Challenge) and LLaMA3-8b (-1.06% on GSM8K, -2.9% on ARC-Easy). As FIPO is a 70b optimizer trained on prompts optimized by GPT-4, its outputs may be too complex for smaller inference models to interpret—indicating a lack of downward compatibility. In contrast, MePO, while built upon a relatively strong base model compared to LLaMA2-7b and LLaMA3-8b Qwen et al. (2025), maintains strong downward compatibility, consistently achieves better performance. More analyses on smaller inference models are in Appx.D.1.
4.3 Ablation Study: Training-Free Comparison
To isolate the effect of learning, we compare MePO with a training-free baseline where the inference model itself rewrites raw prompts using static merit-guided templates (Fig. 14), and then generates responses from the rewritten prompts. This comparison tests whether learning to apply merits is more effective than applying them heuristically.
As shown in Table 3, MePO consistently outperforms the training-free approach across all datasets and model backbones. Notably, average accuracy improves by 2.38% on Tulu2 and over 2.59% on LLaMA2 and 3, demonstrating the value of end-to-end optimization over training-free prompting.
4.4 Case Study: Upward Compatibility
LLaMA2-13b
-chat-hf
Llama-3.3-70B
-Instruct
Tulu2-70b
@@ 4.4 Case Study: Upward Compatibility
- MePO - MePO - MePO
ARC-Easy 49.83 55.93 95.03 92.34 80.47 85.23
ARC-Challenge 45.05 52.73 92.24 91.04 58.87 60.49
GSM8K 28.73 35.97 91.74 92.49 61.41 64.67
BBH 39.49 46.74 66.25 68.25 51.63 53.07
PiQA 56.13 60.99 87.87 89.93 77.75 80.14
Table 4: Response evaluation of MePO prompts in larger LLMs. Detailed BBH results are in Table 9 and 14.
4.5 Case Study: Meta-Prompt Comparison
To further validate MePO’s effectiveness, we compare it on GSM8K with three online-based methods: APO Pryzant et al. (2023), APE Zhou et al. (2022), and PE2 Ye et al. (2024). These methods enhance reasoning by appending an optimized meta-prompt after the query, keeping the original query unchanged while MePO directly optimizes each query. As shown in Table 5, MePO outperforms all baselines, indicating that per-query merit-driven optimization is more effective than applying a fixed meta-prompt across diverse inputs.
Optimizer Tulu2-7b-dpo LLaMA2-7b-chat-hf
APO 29.19 22.97
Iterative APE 29.72 20.92
PE2 32.15 22.37
MePO 35.18 25.25
Table 5: Comparison between meta-prompts and MePO. Details of the applied prompts are in Appx.D.2.
4.6 Case Study: Interpretability of MePO
Example 1
Example 2
These examples illustrate how MePO’s optimized prompts reflect our defined merits—clarity, precision, and concise CoT guidance—to reduce ambiguity and improve task performance.
@@ 5 Related Work
Automatic Prompt Optimization (APO) provides a practical alternative to fine-tuning by optimizing prompts (Yang et al., ; Zhu et al., 2023). Beyond exploration-based methods search over predefined prompt pools (Ma et al., 2023; Shi et al., ; Zhu et al., 2025), recent works favor discrete exploitation-based methods that directly generate or revise prompts without relying on the quality of initial pools, making them more flexible and better suited for generative tasks beyond fixed-answer classification.
Online-Based Discrete APO Several works perform discrete prompt optimization using high-capacity, API-based LLMs as optimization models. For example, Zhou et al. (2022) use InstructGPT; Wang et al. (2024); Guo et al. (2024); Pryzant et al. (2023) rely on GPT-3.5; Xiang et al. (2025) adopt Claude 3.5 Sonnet; and Yuksekgonul et al. (2024) utilize GPT-4o. While effective, these approaches depend on proprietary APIs, introducing practical concerns around cost, inference latency, and privacy risks.
Local Discrete APO Locally trainable prompt optimizers have been explored for general use. FIPO (Lu et al., 2025) trains a Tulu2-based model using a GPT-3.5/4-generated preference dataset. BPO (Cheng et al., 2024) uses a LLaMA-based optimizer trained on ChatGPT-optimal prompts. MAPO (Chen et al., 2023) adopts a model-adaptive strategy with GPT-3.5-generated prompts but requires a separate warm-up set per task, limiting generalization. In contrast, our model uses merit-guided prompts from a lightweight LLM, achieving strong performance without API reliance.
@@ Appendix B Identifying Effective Prompt Merits via Comparative Evaluation
Appendix C POP Dataset Construction and Training Details
@@
Prompt Degradation.
@@
LLaMA3.2-1b-instruct - 27.02 23.13 16.38 31.05 49.56 29.43
Inference Model 28.79 24.23 20.24 32.29 49.73 31.06
BPO 27.99 24.04 15.85 31.77 49.89 29.91
FIPO 25.19 23.55 21.30 31.72 50.27 30.41
MePO 29.42 24.40 21.46 33.40 50.44 31.82
LLaMA3.2-3b-instruct - 73.32 55.46 63.23 35.60 66.32 58.79
Inference Model 78.16 60.75 63.00 37.91 69.37 61.84
BPO 74.75 55.74 56.71 37.34 67.14 58.34
FIPO 78.62 60.98 65.13 37.65 69.53 62.38
MePO 79.55 61.43 65.66 40.49 70.29 63.48
@@ Appendix D Dataset Information and Experimental Implementation
We evaluate MePO on four multiple-choice and one generative benchmark dataset:
@@
We compare MePO with two state-of-the-art prompt optimizers. The prompt instructions used for inference models serving as prompt optimizers are shown in Fig. 11. Details of the baselines are as follows:
@@
D.1 Case Study: Downward Compatibility
To further analyze the downward compatibility of optimization models, we evaluate performance using two smaller inference models: LLaMA-3.2-3B-Instruct [^13] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct and LLaMA-3.2-1B-Instruct [^14] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct.
As shown in Table 7, both BPO and FIPO exhibit performance degradation when paired with lightweight inference models. For instance, under LLaMA-3.2-3B-Instruct, compared to using the inference model itself as the optimizer, BPO drops by 3.41% on ARC-Easy, 5.01% on ARC-Challenge, 6.29% on GSM8K, and 2.23% on PiQA. Under LLaMA-3.2-1B-Instruct, FIPO drops by 3.60% on ARC-Easy, and BPO drops by 4.39% on GSM8K. Both also show decreased performance on BBH across the two lightweight inference models. In contrast, MePO consistently improves performance across all datasets, demonstrating strong downward compatibility and robustness in low-resource settings.
D.2 Case Study: Online-Based Discrete APO Comparison
To further validate MePO’s effectiveness in optimizing prompts or instructions with refined meta-prompts, we compare MePO with three discrete APO methods on GSM8K:
  • APO, a prompt optimization method that performs optimization via textual gradients using GPT-3.5-Turbo-Instruct in a text-based dialogue setting;
  • Iterative APE, a template-based strategy in which GPT-3.5 generates a pool of candidate prompts, followed by selection based on development set performance;
  • PE2, a prompt engineering method that optimizes prompts through three key components using GPT-3.5-Turbo-Instruct.
Optimizer Meta Prompt
APO
Given the scenario, perform necessary
calculations and provide a step-by-step
explanation to arrive at the correct
numerical answer. Consider all
information provided.
Iterative APE
Let’s dissect this and tackle it gradually,
one phase at a time.
PE2
Let’s solve the problem step-by-step and
calculate the required total value correctly.
D.3 BBH Experimental Results
Detailed results for the 23 BBH tasks are presented in the following tables: Table 9 (raw prompts), Table 10 (optimized by the inference model), Table 11 (training-free merit-guided optimization), Table 12 (BPO), Table 13 (FIPO), and Table 14 (MePO).
- Qwen2-7b
@@ D.3 BBH Experimental Results
date_understanding 44.8 30.8 28.8 31.6 42 25.6 27.2 62 60.8
disambiguation_qa 63.2 32.8 31.2 32.4 30 36.4 33.6 68.8 65.6
hyperbaton 75.6 48.4 48.4 48.4 48.4 51.6 48.4 78 87.6
logical_deduction_five_objects 49.2 22.4 18 28 33.2 24.4 19.6 50 58.8
logical_deduction_seven_objects 42.4 17.6 14.8 30 27.2 30.4 15.2 42.8 54.4
logical_deduction_three_objects 66.4 32.8 31.6 33.6 36.4 35.2 31.6 71.2 75.6
movie_recommendation 48 26.8 22.4 33.2 23.6 23.2 22.4 52.8 76.4
penguins_in_a_table 54.11 21.23 22.6 31.51 46.58 26.71 23.97 50.68 67.12
reasoning_about_colored_objects 47.6 32.4 32 16.4 29.2 30 16.4 46 51.6
ruin_names 37.2 24 30.8 35.2 27.2 41.6 28.4 66.8 51.2
salient_translation_error_detection 32 26.8 23.2 21.6 28 34 16.8 52 71.6
snarks 60.11 44.94 46.07 46.07 50.56 57.3 46.07 81.46 74.16
word_sorting 88 82.4 80 84 80 52 1.2 100 100
temporal_sequences 61.2 19.2 28.4 28.4 33.2 28 28.4 62 83.6
tracking_shuffled_objects_five_objects 28.4 19.2 20 20.8 30.8 20 20 17.2 40.4
tracking_shuffled_objects_seven_objects 13.2 24 14.4 13.6 22 11.2 12.8 13.6 18.4
tracking_shuffled_objects_three_objects 27.2 29.2 24.4 26.8 49.2 30 30.8 28 31.2
causal_judgement 62.57 49.2 51.87 51.34 49.2 47.59 51.87 53.48 79.68
formal_fallacies 54 52.8 53.2 53.6 52.4 51.6 53.2 52 77.2
navigate 57.6 53.6 42 41.6 41.6 41.2 42 52.8 57.2
web_of_lies 46.4 48.8 48.8 49.2 54 44.4 46.4 46.8 68.4
sports_understanding 70 50 46.4 70.4 54 53.6 43.2 45.6 66.8
boolean_expressions 34.8 56.8 52.4 54.8 51.2 48 54 30.4 62.8
multistep_arithmetic_two 33.6 12.4 9.6 36 14 46 16 16.8 85.6
object_counting 36 33.2 32.8 50.4 33.2 48.8 46.8 49.6 90
@@
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
date_understanding 54.4 34 34 32 27.6 31.2
disambiguation_qa 62.8 34 32.4 34.4 40.4 36.8
hyperbaton 77.6 49.2 49.2 49.2 51.6 48.4
logical_deduction_five_objects 50.4 22.8 19.2 36.4 29.2 20.8
logical_deduction_seven_objects 44 17.2 15.6 31.2 29.6 16.4
logical_deduction_three_objects 74.8 34.8 32 38.8 40 32
movie_recommendation 51.6 30.4 23.6 35.2 24.4 26.8
penguins_in_a_table 56.16 26.03 23.29 32.19 28.77 23.29
reasoning_about_colored_objects 48.4 35.2 34 18.8 28.4 18
ruin_names 38.4 26.4 32.4 38 43.6 28.4
salient_translation_error_detection 38 21.2 24.8 23.6 35.2 20.4
snarks 65.73 48.88 49.44 47.19 57.3 46.07
word_sorting 90.8 84 84 86.4 60 1.2
temporal_sequences 62.8 31.2 33.2 30.4 28.4 28.4
tracking_shuffled_objects_five_objects 29.6 29.6 30.8 32.4 20 22.4
tracking_shuffled_objects_seven_objects 15.2 25.6 23.2 14.4 14 13.6
tracking_shuffled_objects_three_objects 28 30 26 28.4 31.2 32
causal_judgement 72.73 52.41 53.2 50.8 55.61 52.41
formal_fallacies 54.8 54.4 53.2 51.6 52 53.2
navigate 60.8 56 43.2 42.8 41.6 42
web_of_lies 47.2 48.8 48.8 50 48.8 47.6
sports_understanding 75.6 63.45 48 72 57.6 45.6
boolean_expressions 41.6 57.2 53.6 52.4 47.6 54
multistep_arithmetic_two 36.4 10.8 12 37.2 54 18
object_counting 39.2 33.2 37.2 58.4 56 48.4
- Qwen2-7b Tulu2-7b-dpo LLaMA2-7b-chat-hf LLaMA3-8b-instruct
date_understanding 55.6 36 34 48
disambiguation_qa 66 37.2 35.2 35.2
hyperbaton 78.4 50 48.4 49.6
logical_deduction_five_objects 55.6 25.2 19.2 34.8
logical_deduction_seven_objects 44 28.4 21.6 38.8
logical_deduction_three_objects 75.2 34.4 42.4 40.4
movie_recommendation 53.2 31.2 24.4 37.6
penguins_in_a_table 56.16 27.4 23.29 36.99
reasoning_about_colored_objects 50.8 34 35.2 20.8
ruin_names 38 28.4 40 39.6
salient_translation_error_detection 38.4 34 37.2 18.8
snarks 62.36 50 50.56 48.88
word_sorting 94.4 87.2 87.6 90
temporal_sequences 64.4 36.8 36.8 31.2
tracking_shuffled_objects_five_objects 29.6 38.4 34.8 33.2
tracking_shuffled_objects_seven_objects 14.4 24.8 23.2 14.4
tracking_shuffled_objects_three_objects 30 31.6 31.6 31.2
causal_judgement 55.08 52.41 51.87 52.41
formal_fallacies 54 54 54 52
navigate 57.2 58.8 43.2 42.4
web_of_lies 52.4 49.6 49.2 51.6
sports_understanding 76 64.8 49.2 79.6
boolean_expressions 41.6 60.4 53.6 50
multistep_arithmetic_two 41.2 12.6 12 41.6
object_counting 39.2 36.4 58.4 50.8
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
date_understanding 56.8 33.2 35.2 48.8 29.6 29.2
disambiguation_qa 63.2 34 34.4 40.4 44.4 34.4
hyperbaton 80 53.2 50 49.6 51.6 48.8
logical_deduction_five_objects 52.8 25.2 24 31.2 29.2 20.4
logical_deduction_seven_objects 46.4 28.8 19.2 35.6 24.8 15.2
logical_deduction_three_objects 78.4 32.8 45.6 39.2 34.8 32
movie_recommendation 61.2 22.8 24.4 37.6 23.2 26.8
penguins_in_a_table 54.11 23.97 23.29 36.3 29.45 23.97
reasoning_about_colored_objects 52.4 34 35.6 20.8 28.4 18
ruin_names 38.4 28 36.8 35.6 38 28.4
salient_translation_error_detection 38.4 27.6 35.2 21.2 33.6 18.8
snarks 66.85 53.37 53.37 48.88 53.93 46.07
word_sorting 92.4 90 87.6 90 72 3.2
temporal_sequences 66.4 26.4 38.4 30.8 35.6 28.4
tracking_shuffled_objects_five_objects 29.6 29.2 31.2 34 20 19.6
tracking_shuffled_objects_seven_objects 14.4 28 23.2 15.2 14 13.2
tracking_shuffled_objects_three_objects 27.6 31.6 31.6 31.6 31.6 32
causal_judgement 53.48 50.8 52.04 52.94 52.94 51.87
formal_fallacies 52 53.2 54.14 54 52.4 53.2
navigate 59.6 58.4 44.8 48.4 44.8 42
web_of_lies 49.2 51.6 48.8 49.6 49.2 48.8
sports_understanding 76.4 61.2 53.2 74.8 55.2 45.6
boolean_expressions 45.6 63.6 55.6 53.2 50.4 54
multistep_arithmetic_two 36.8 10 17.6 37.2 46.8 16.8
object_counting 34.4 32.4 34.8 50 54.4 43.6
- Qwen2-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
date_understanding 55.2 38.4 29.2 36 30 27.2
disambiguation_qa 61.6 34.8 31.2 40.4 32 33.2
hyperbaton 78.8 33.2 48.4 60 52 48.4
logical_deduction_five_objects 50 21.6 22.8 34.4 23.6 20.4
logical_deduction_seven_objects 46.4 25.2 20.4 39.2 23.2 15.6
logical_deduction_three_objects 70.8 30.4 42 39.2 38.8 32
movie_recommendation 62 36.8 23.97 31.6 34.4 24.4
penguins_in_a_table 54.11 36.3 23.29 30.82 52.05 23.29
reasoning_about_colored_objects 50.8 31.6 33.2 24.8 25.6 16.4
ruin_names 46 34.4 36 38.8 40 29.2
salient_translation_error_detection 37.6 32.4 33.6 24.4 37.2 18.8
snarks 62.92 46.07 50.56 46.07 55.62 46.07
word_sorting 92.4 92.4 86.4 92 70 4.8
temporal_sequences 63.2 27.6 28.4 28.8 40 28.8
tracking_shuffled_objects_five_objects 31.2 24.8 32.4 36.8 20.8 20.8
tracking_shuffled_objects_seven_objects 16.8 25.6 24.8 14.4 14.4 16.4
tracking_shuffled_objects_three_objects 31.2 30.4 31.6 30.4 31.6 30.8
causal_judgement 62.57 53.48 51.87 51.34 55.08 52.41
formal_fallacies 53.2 54.4 54 52.4 53.2 53.2
navigate 59.2 60.8 42 46.4 53.2 42
web_of_lies 58.8 48 53.2 45.2 49.2 48.8
sports_understanding 81.2 66.4 47.26 71.6 55.6 46
boolean_expressions 39.6 58.8 53.6 61.2 39.6 54.4
multistep_arithmetic_two 16.4 11.6 17.2 39.6 24.4 16.4
object_counting 36.4 33.6 61.2 50.8 55.2 43.2
- Qwen2-7b

### v3

@@ 4.4 Ablation Study: Training Ablation
MePO 83.33 68.52 83.12 54.35 83.46 74.56
Tulu2-7b
@@
Training Free 49.58 34.30 32.93 40.96 72.60 46.22
MePO 55.05 38.14 35.18 43.25 73.60 49.04
LLaMA2-
7b-chat-hf
Training Free 36.91 30.55 25.23 39.88 53.10 37.13
MePO 39.86 31.74 29.42 41.97 55.33 39.66
Gemma2-9b Training Free 91.33 48.32 67.17 66.23 85.00 71.61
MePO 92.30 48.89 68.67 69.47 85.35 72.94
Qwen3-8b Training Free 83.63 54.57 89.84 70.82 82.54 76.28
MePO 84.38 54.68 91.12 71.12 83.08 76.88
Table 4: Comparison between training-free prompts and MePO. Detailed BBH results are in Appx.D.1.
To isolate the effect of learning, we compare MePO with a training-free baseline where the inference model itself rewrites raw prompts using static merit-guided templates (Fig. 14), and then generates responses from the rewritten prompts. This comparison tests whether learning to apply merits is more effective than applying them heuristically. As shown in Table 4, MePO outperforms the training-free approach across all datasets and model backbones. The average accuracy improves by 1.01%, 2.38%, 2.53%, 1.33%, and 0.6% on Qwen2, Tulu2, LLaMA2, Gemma2 and Qwen3, demonstrating the value of end-to-end optimization over training-free prompting.
4.5 Ablation Study: Merit Ablation
To assess the contribution of each merit proposed in Sec. 2, we conduct an ablation study targeting four designed merits under the training-free setup: (1) Clarity, (2) Precision, (3) Concise CoT, and (4) Preservation. We evaluate on ARC-Easy, ARC-Challenge, GSM8K, and PiQA under Qwen2-7B and Vicuna under Qwen-7B-Chat.
[…]
- 80.68 65.19 76.88 80.20 75.74 -
MePO 82.79 68.17 81.52 82.32 78.70 33.5
MePO_wo1 80.01 66.21 77.58 80.52 76.08 12.5
MePO_wo2 82.03 68.00 77.03 80.25 76.83 15.0
MePO_wo3 81.31 67.32 81.27 81.18 77.77 21.3
MePO_wo4 81.48 67.66 80.82 81.50 77.87 27.5
Table 5: Merit Ablation. MePO_wo1–MePO_wo4 each remove one merit while keeping the others. All variants follow the template in Fig.14, excluding the targeted merit. “-” corresponds to the original dataset prompts without any optimization. In Vicuna comparison, we evaluate the $\Delta$ Win Rate between MePO variants and without optimization.
As shown in Table 5, all MePO variants outperform the raw prompts, demonstrating the overall effectiveness of merit-driven optimization. The removal of Clarity (MePO_wo1) and Precision (MePO_wo2) results in the most significant performance drops, consistent with Fig.2, where these merits were most frequently associated with advanced LLMs-identified “good prompts”.
Notably, removing Concise CoT (MePO_wo3) has minimal impact, even on GSM8K, suggesting that when prompts are already clear and precise, LLMs can reason effectively without explicit CoT prompting. This also helps explain why LLMs sometimes fail on raw prompts but succeed with CoT: the benefit of CoT may stem not from added reasoning capacity but from clarifying vague inputs. When clarity and precision are provided directly, CoT prompting becomes less necessary.
4.6 Further Analyses
More in-depth analyses are in Appx.E, including downward compatibility with weaker LLMs (E.1), upward compatibility with larger LLMs (E.2), newly released LLM comparison (E.3), interpretability and examples of MePO (E.4), dataset ablation studies (E.5), POIR comparison (E.6), significant test (E.7), online-based meta-prompts comparison (E.8); and iterative optimization (E.9).
@@ 5 Related Work
Automatic Prompt Optimization (APO) provides a practical alternative to fine-tuning by optimizing prompts (Yang et al., ; Zhu et al., 2023). Beyond exploration-based methods search over predefined prompt pools (Ma et al., 2023; Shi et al., ; Zhu et al., 2025, 2024), recent works favor discrete exploitation-based methods that directly generate or revise prompts without relying on the quality of initial pools, making them more flexible and better suited for generative tasks beyond fixed-answer classification.
Online-Based Discrete APO Several works perform discrete prompt optimization using high-capacity, API-based LLMs as optimization models. For example, Zhou et al. (2022) use InstructGPT; Wang et al. (2024); Guo et al. (2024); Pryzant et al. (2023) rely on GPT-3.5; Xiang et al. (2025) adopt Claude 3.5; and Yuksekgonul et al. (2024) utilize GPT-4o. While effective, these approaches depend on proprietary APIs, introducing practical concerns around cost, inference latency, and privacy risks.
Local Discrete APO Locally trainable prompt optimizers have been explored for general use. FIPO (Lu et al., 2025) trains a Tulu2-based model using a GPT3.5/4-generated dataset. BPO (Cheng et al., 2024) uses a LLaMA-based optimizer trained on ChatGPT-optimal prompts. POIR Liang et al. (2024) implicitly trains a lightweight PO using model self-reasoning and self-evaluation. MAPO (Chen et al., 2023) adopts a model-adaptive strategy with GPT-3.5-generated prompts but is task-specific, limiting generalization. In contrast, our model uses explicit merit-guided prompts from a lightweight LLM to enable a high-performance, locally deployable task-agnostic optimizer.
@@ Appendix B Identifying Effective Prompt Merits via Comparative Evaluation
Appendix C POP Dataset Construction, Evaluation and Training Details
@@
C.1 Human Annotator Evaluation Results.
Four human annotators evaluated 200 samples optimized from each of the Alpaca and BPO datasets in our constructed POP dataset, assessing both prompt-level and response-level quality. The results are shown in Table 6.
Alpaca BPO
Prompt Win Rate
MePO
Win
Tie
Raw
Win
MePO
Win
Tie
Raw
Win
HE1 38.9% 41.9% 19.2% 75.4% 21.6% 3.0%
HE2 71.8% 12.6% 15.6% 71.2% 16.2% 12.6%
HE3 88.0% 9.0% 3.0% 80.9% 9.6% 9.5%
HE4 59.6% 18.7% 21.7% 58.3% 29.1% 12.6%
Response Win Rate
MePO
Win
Tie
Raw
Win
MePO
Win
Tie
Raw
Win
HE1 54.8% 29.1% 16.1% 59.6% 9.1% 31.3%
HE2 65.8% 12.6% 21.6% 84.3% 12.6% 3.0%
HE3 77.8% 12.6% 9.6% 75.0% 12.5% 12.5%
HE4 62.8% 15.6% 21.6% 71.2% 19.2% 9.6%
Table 6: Human evaluation results on the POP dataset. Each human evaluator (HEi) compares MePO-optimized prompts against raw prompts on samples from the Alpaca and BPO datasets.
We also ask the four annotators to select one or more merits for each preferred prompt and report their preferences based on the selected merits. The merit recognition results are presented in E.5.2.
C.2 Dataset Evaluation API Cost.
In our method, the primary cost arises from dataset evaluation and filtering using DeepSeek-R1 to assess prompt and response quality for constructing the (P_silver, P_golden, R_silver, R_golden) tuples. The total DeepSeek-R1 token usage was: Input Cache Hit: 55,060,029; Input Cache Miss: 53,429,910; and Output Tokens: 29,892,599. The corresponding cost is approximately 40.67 USD. This usage occurred between April 8 and May 1, 2025. Notably, DeepSeek offered off-peak discounts during this period, which may have affected the final token billing.
C.3 Dataset Under-performance Case Analysis.
Although MePO prompts were preferred in most cases, a small portion was disfavored during dataset construction. We conducted an error analysis from both human evaluation and DeepSeek evaluation perspectives to understand the underlying reasons.
From the human evaluation side, annotators occasionally favored simpler raw prompts, even when MePO versions offered greater specificity.
BPO Example.
Raw Prompt (Preferred): If someone is having second thoughts about something, what should I tell him/her to calm down?
MePO Prompt: When someone is experiencing second thoughts about a decision, provide reassurance by emphasizing the following points clearly and concisely: 1. Reiterate the positive aspects and rationale behind the initial decision. 2. Highlight past experiences where similar concerns were resolved positively. 3. Encourage reflective thinking to identify specific reasons for the current doubt and consider their validity. 4. Suggest taking a brief break to regain perspective before making a final judgment. 5. Remind them of their own capabilities and past successes in handling similar situations. 6. Offer emotional support and validation of their feelings without dismissing them. By guiding them through these steps, help them regain confidence and make an informed decision calmly.
Alpaca Example.
Raw Prompt (Preferred): Compare the differences between a wooden and plastic baseball bat.
MePO Prompt: Identify and describe with precision the key differences in design, material properties, performance characteristics, and player preferences between a traditional wooden baseball bat and a modern plastic baseball bat, providing a detailed comparative analysis.
These examples suggest a preference among some annotators for brevity and clarity over elaboration, particularly in straightforward tasks.
From the DeepSeek evaluation side, preference differences arose from subtle distinctions in precision and phrasing naturalness. Representative comments include:
Comment 1. “Both prompts are very similar, but ‘force on a 1 kg mass due to the gravitational force’ is slightly more precise because it explicitly mentions the cause of the force (gravitational force), which could help avoid any ambiguity.”
Comment 2. “##Prompt 2## is slightly better because it is more concise and direct, while still clearly conveying the task. The instruction ‘Rewrite the following sentence in a more formal way’ is straightforward and leaves no ambiguity. ##Prompt 1## is also clear but slightly more verbose without adding value.”
Comment 3. “Prompt 2 is more concise and direct, which is generally better for eliciting a straightforward answer. Prompt 1’s request for a ‘step-by-step explanation’ is unnecessary for a simple fact like this, and may lead to overly verbose responses.”
In sum, these cases highlight that optimized prompts, while generally preferred, may underperform when verbosity outweighs clarity or when evaluators favor conciseness in simple tasks.
C.4 Prompt Degradation.
@@ Appendix D Dataset Information and Experimental Implementation
We evaluate MePO on question-answering, math, and instruction-following benchmark datasets:
@@
We compare MePO with two state-of-the-art prompt optimizers. The prompt instructions used for inference models (Qwen2-7b [^15] https://huggingface.co/Qwen/Qwen2-7B, Tulu2-7b-dpo [^16] https://huggingface.co/allenai/tulu-2-dpo-7b, LLaMA2-7b-chat-hf [^17] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Gemma2-9b [^18] https://huggingface.co/google/gemma-2-9b, and Qwen3-8b [^19] https://huggingface.co/Qwen/Qwen3-8B) serving as prompt optimizers are shown in Fig. 11. Details of the baselines are as follows:
@@
D.1 BBH Experimental Results
Detailed results for the 25 BBH tasks are presented in the following tables: Table 8 (raw prompts), Table 9 (optimized by the inference model), Table 10 (training-free merit-guided optimization), Table 11 (BPO), Table 12 (FIPO), and Table 13 (MePO).
Qwen2
-7b
@@ D.1 BBH Experimental Results
Gemma2
-9b
date_understanding 44.8 30.8 28.8 31.6 42 25.6 27.2 62 60.8 51.6
disambiguation_qa 63.2 32.8 31.2 32.4 30 36.4 33.6 68.8 65.6 67.6
hyperbaton 75.6 48.4 48.4 48.4 48.4 51.6 48.4 78 87.6 85.2
logical_deduction_five_objects 49.2 22.4 18 28 33.2 24.4 19.6 50 58.8 62.8
logical_deduction_seven_objects 42.4 17.6 14.8 30 27.2 30.4 15.2 42.8 54.4 65
logical_deduction_three_objects 66.4 32.8 31.6 33.6 36.4 35.2 31.6 71.2 75.6 51.6
movie_recommendation 48 26.8 22.4 33.2 23.6 23.2 22.4 52.8 76.4 66.4
penguins_in_a_table 54.11 21.23 22.6 31.51 46.58 26.71 23.97 50.68 67.12 57.53
reasoning_about_colored_objects 47.6 32.4 32 16.4 29.2 30 16.4 46 51.6 30
ruin_names 37.2 24 30.8 35.2 27.2 41.6 28.4 66.8 51.2 45.6
salient_translation_error_detection 32 26.8 23.2 21.6 28 34 16.8 52 71.6 66.4
snarks 60.11 44.94 46.07 46.07 50.56 57.3 46.07 81.46 74.16 67.56
word_sorting 88 82.4 80 84 80 52 1.2 100 100 90
temporal_sequences 61.2 19.2 28.4 28.4 33.2 28 28.4 62 83.6 52.8
tracking_shuffled_objects_five_objects 28.4 19.2 20 20.8 30.8 20 20 17.2 40.4 30
tracking_shuffled_objects_seven_objects 13.2 24 14.4 13.6 22 11.2 12.8 13.6 18.4 28.8
tracking_shuffled_objects_three_objects 27.2 29.2 24.4 26.8 49.2 30 30.8 28 31.2 45.6
causal_judgement 62.57 49.2 51.87 51.34 49.2 47.59 51.87 53.48 79.68 87.17
formal_fallacies 54 52.8 53.2 53.6 52.4 51.6 53.2 52 77.2 54.4
navigate 57.6 53.6 42 41.6 41.6 41.2 42 52.8 57.2 81.2
web_of_lies 46.4 48.8 48.8 49.2 54 44.4 46.4 46.8 68.4 68.4
sports_understanding 70 50 46.4 70.4 54 53.6 43.2 45.6 66.8 68.4
boolean_expressions 34.8 56.8 52.4 54.8 51.2 48 54 30.4 62.8 64.8
multistep_arithmetic_two 33.6 12.4 9.6 36 14 46 16 16.8 85.6 75.2
object_counting 36 33.2 32.8 50.4 33.2 48.8 46.8 49.6 90 71.2
@@
Qwen2
-7b